Fermi National Laboratory
 

Operation Summary for February 9, 2011 - February 11, 2011

 

Wednesday Feb. 9

The day shift began with the Tevatron (TEV) in shot setup, with Operations monitoring the Recycler stash (400.6E10), with the Antiproton Source (Pbar) stacking (21.1E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Pbar personnel accessed the AP0 water cage. Operations established store 8486 at 9 AM with an average initial luminosity (AIL) of 327E30. MI personnel conducted studies during the afternoon.

During the evening shift, the MI studies concluded early on the evening shift. Operations reported that a Linac RF station (LRF1) tripped off numerous times. Linac personnel investigated and returned the station to service.

Thursday Feb. 10

The midnight shift began with Operations monitoring store 8486 (80.73E30) and the stash (343.6E10), with Pbar stacking (25.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the MI power supply interlock loop tripped off at 5:35 AM on a ground fault indication. Operators isolated the ground fault to the Lower bus between MI-30 and MI-40. Operators had MI prepared for access by the end of the shift. Linac personnel began work on a RF station (LRF2) at 7:40 AM.

During the day shift, MI personnel accessed to search for the ground fault, which they discovered was due to a cable tray grounding strap that came lose from the wall and fell against the MI Lower bend bus. MI personnel reattached the strap to the wall and came out of access. Operations resumed sending beam to MiniBooNE at 1:35 PM. Operations resumed sending beam to Pbar and NuMI at 2:35 PM. TEV personnel began conducting studies.

During the evening shift, Operations terminated store 8486 at 7:10 PM, put the TEV into HEP shot setup, and established store 8487 at 10:30 PM with an AIL of 162.72E30.

Friday Feb. 11

The midnight shift began with Operations monitoring store 8487 (93.97E30) and the stash (113.5E10), with Pbar stacking (22.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that a Booster RF station’s (BRF13) gradient began to weaken; operators turned off BRF13 and brought BRF19 online in its place.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct slipstacking and gap clearing studies and allow Booster RF personnel to work on BRF13.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/11/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory